# Rambo: Enhancing RAG-based Repository-Level Method Body Completion

### v1

@@ 1 Introduction
We conducted several experiments to evaluate the performance of Rambo in repo-level method body completion using state-of-the-art code LLMs, including DeepSeek Coder [7], Code Llama [6], and GPT-3.5 Turbo [19]. These experiments were conducted across 40 Java projects with a total of 674 repo-level MBC problems. Our results demonstrate that Rambo significantly surpasses RepoCoder across all evaluated code LLMs and metrics. Specifically, Rambo achieves impressive improvements over the other MBC approaches in various performance metrics, by up to 46% in BLEU, up to 57% in CodeBLEU, up to 36% in Compilation Rate, and up to 3X in Exact Match rate. Remarkably, Rambo delivers up to 12% enhancement in the Exact Match rate compared to the Oracle RAG-based method [15], which estimates the upper-bound of performance for RepoCoder and other methods employing the same pipeline. This substantial improvement underscores Rambo’s superior capability in generating high-quality method bodies and sets a new benchmark in the task of repository-level method body completion.
@@
2. Advanced Retrieval Strategy for RAG-based MBC: We introduce a sophisticated retrieval mechanism that identifies and utilizes essential repo-specific code elements, such as classes, methods, and fields, to ensure more accurate and contextually relevant method body generation.
@@ 3.1 Repository-specific Context Retrieval
This section describes the retrieval phase, $C_{R}=\textit{retrieve}(m,R)$, in Rambo to search for the repo-specific context relevant to the given infilling method. Our retrieval approach consists of two phases, Essential Code Element Identification and Relevant Usage Extraction.
@@ 3.1.1 Essential Code Element Identification (EEI)
To enhance the relevance of the sketch, Rambo first searches for the set of methods, $L$, within the repository $R$ whose interface (input-output) and name are most similar to those of $m$. The set of methods $L$ serves as the reference points for understanding the expected structure and functionality of $m$. After that, Rambo employs a RAG-based method to generate a sketch code $\mathcal{s}$ of the method body using the retrieved methods $L$. Although the sketch, generated as $\mathcal{s}=\mathscr{M}(m,L)$, may not fully accurate, $\mathcal{s}$ includes potentially essential code elements such as called methods, accessed fields, and used types. These elements are then analyzed and extracted:
@@
where $\mathcal{M}_{m}$, $\mathcal{M}_{f}$, and $\mathcal{M}_{t}$ are the sets of called methods, accessed fields, and used types in the sketch $\mathcal{s}$, respectively. Note that analyzing the generated sketch code $\mathcal{s}$ by existing code analyzers (e.g., JDT or Joern [20]) is feasible because the code generated by pre-trained code models, especially code LLMs, is quite syntactically correct [21, 13]. However, the employed code LLM may hallucinate code, possibly generate the code elements not existing/declared in the repo or use them inappropriately [14, 13].
To mitigate code LLM hallucination, after analyzing, Rambo searches for similar code elements already declared within the repository. This involves identifying the sets of methods, fields, and classes whose names are most similar to those found after analyzing $\mathcal{s}$:
@@
$E_{c}=\bigcup_{\begin{subarray}{c}e\in\mathcal{M}_{t}\end{subarray}}\left\{\arg\max\limits_{c\in\text{cls}(R)}\left(\text{name\_sim}(e,c)\right)\right\}$ (3)
@@ 3.1.2 Relevant Usage Extraction (RUE)
Once these usages are identified, they are ranked based on their relevancy to the functionality of the method $m$. Particularly, the relevancy of a usage $u$ is calculated by function rel:
@@
where $\text{sim}(u,\mathcal{s})$ calculates the similarity between a usage $u\in U$ and the sketch $\mathcal{s}$ generated during EEI. The reason for this estimation of $\text{rel}(u)$ is that the body of $u$ expresses how an essential element $e$ is utilized and $\mathcal{s}$ roughly expresses how potentially essential elements should be used. Thus, if the usage of $e$ is also present in the sketch code, it indicates that the usage of $e$ in $u$ is likely relevant to $m$. This ranking ensures that the most relevant usages are prioritized, providing the code LLM with high-quality context that is both specific to the repository and relevant to the task at hand. Additionally, Rambo brings back the set of methods $L$ whose signature closely matches that of $m$, serving as a structural and functional reference point for method $m$. The methods in $L$ and the top relevant usages are concatenated to form the repo-specific relevant context $C_{R}$. Finally, this context and the surrounding context $C_{f}$ of $m$ are fed into $\mathscr{M}$ to generate the final method body $\widehat{m_{b}}=\mathscr{M}(m,C_{f},C_{R})$.
@@ 3.2 Method Body Generation
To generate $m$’s body, we design a prompt template that seamlessly integrates the identified relevant context $C_{R}$ with the infilling method, embedding the context as code comments within the class (as illustrated in Fig. 1). The retrieved usages are arranged in ascending order of relevancy, with each code snippet annotated by its original file path. The number of snippets included in the prompt is optimized based on the available prompt length. This approach ensures that the prompt is packed with relevant information, facilitating the generation of a method body that is both functionally accurate and consistent with the repository’s coding standards.
@@ 4.1 Dataset
In this work, we evaluate the MBC performance of Rambo and other approaches on two large MBC benchmarks (Table I). First, we utilize Defects4J [22] as in CoderUJB [23], a multi-task benchmark for several SE tasks such as functional code generation, test generation, defect detection, and automated program repair. In this work, we adapt the functional code generation task from Defects4J in CoderUJB to the MBC task, resulting in 204 MBC problems in 14 real open-source Java projects. Additionally, to evaluate the performance of the repo-level MBC approaches on more up-to-date repositories, we construct another benchmark which includes active high-quality open-source repositories from GitHub. Specifically, we collect 26 Java repositories (other than those in Defects4J) that meet the following criteria: they must have an open-source license, have been actively updated, possess over 2K stars, and be compilable. From each non-test class in the collected projects, we randomly select a method to generate an MBC problem and the corresponding reference, which is the method’s body. To ensure a reasonable difficulty level of the benchmark, we exclude all constructors, getters, setters, and too-short methods with fewer than 3 LOCs. As a result, we obtain a total of 470 MBC problems. In this work, for every MBC problem, we omitted the import section in the left context, $C_{f}^{\text{left}}$, to accurately reflect the coding activities in practice, that the necessary imports are usually unavailable before implementing the infilling method.
TABLE I: Datasets statistics
@@ 4.2.1 Empirical Procedure
Baselines. We compared Rambo against the state-of-the-art Repo-level Method Body Completion (MBC) approaches:
@@
    RepoCoder [15]: RepoCoder employs a similarity-based retriever and a code LLM in an iterative retrieval-generation pipeline.
@@
    RepoCoder Oracle [15]: This method estimates the upper bound of RepoCoder’s performance by utilizing the ground truth code to search for relevant context in the RepoCoder. Note that this method performs a single retrieval process to obtain relevant code snippets.
@@
    Rambo Oracle: Inspired by RepoCoder Oracle [15], we estimated the upper bound of Rambo’s performance by its own Oracle. Within Rambo’s pipeline, this oracle method directly considers the ground truth code as the sketch code $\mathcal{s}$ in EEI and REU. In this case, the essential elements are correctly identified, and their usages are ranked based on the similarities with the ground truth.
We used the implementation in their original papers for all the baseline approaches. For RepoCoder and RepoCoder Oracle, we used the hyper-parameters recommended in the original work [15]. Additionally, we applied RepoCoder with a single iteration at which RepoCoder achieves the best overall performance with our studied dataset. We evaluated the MBC performance of the approaches under study using several representative code LLMs, both open-source and commercial, including DeepSeek Coder 6.7B [7], Code Llama 7B [6], and GPT-3.5 Turbo [19] with about 175B parameters. For detailed implementation and other hyper-parameters, one can see our website [26]. All experiments were run on a machine with an Intel Xeon Bronze 3104 (6 cores) @ 1.70GHz, 32GB RAM, and an RTX A6000 GPU.
@@
  • Embedding Techniques: We evaluated how different embedding methods influence the retrieval of code elements, relevant usages, and MBC performance.
  • Relevant Context Size: We analyzed how changes in the relevant contexts’ size influence the MBC performance.
  • Code LLM’s Size: We studied the effect of code LLM sizes on Rambo’s performance, balancing computational efficiency and the quality of the generated code.
@@
  • Repository Size: We examined the impact of repository sizes on Rambo’s ability to retrieve relevant context and generate functionally accurate method bodies.
@@ 4.2.2 Metrics
  • BLEU [27] is a widely used metric in NLP, assessing the similarity between generated and reference text by comparing overlapping n-grams. In MBC, it measures surface-level similarity, focusing on the direct overlap of tokens between the generated and reference code.
  • CodeBLEU [28] extends the BLEU score by incorporating code-specific properties, such as syntax and semantic structure. It measures the alignment between the generated and the reference codes, capturing surface-level similarities and deeper structural elements.
  • Compilation Rate (CR) assesses the percentage of generated code snippets that successfully compile without errors. A higher compilation rate reflects the syntactical correctness of the generated code and its adherence to the language’s rules, making it a crucial metric for evaluating the practicality of code generation models.
  • pass@k [29] evaluates the generated code’s ability to pass a set of predefined unit tests, measuring the percentage of solutions that pass all the given tests when considering the top-k predictions. A higher pass@k value indicates stronger model performance in producing functionally correct code. In this work, $k=1$ and pass@k is applicable for evaluation using Defects4J.
@@ 5.1 Answering RQ1: Accuracy and Comparison
TABLE II: Repo-level MBC performance comparison
Rambo’s Benchmark Defects4J
Code LLM BLEU CodeBLEU CR EM BLEU CodeBLEU CR pass@k EM
CodeLlama (7B) Rambo Oracle 41.83 46.22 51.28 14.89 54.19 55.90 64.71 46.08 19.12
RepoCoder Oracle 46.59 49.53 55.96 14.68 53.69 54.47 64.71 46.57 19.61
In-File 25.09 33.85 48.09 3.83 44.06 46.73 55.88 33.82 8.82
RepoCoder 42.98 47.66 55.32 12.77 51.27 53.46 60.78 46.57 20.10
Rambo 43.16 47.76 55.96 13.62 52.04 54.98 63.73 44.12 20.59
DeepSeek Coder (6.7B) Rambo Oracle 50.87 53.96 63.40 18.72 51.31 53.54 67.16 49.51 21.57
RepoCoder Oracle 50.04 52.48 58.51 17.02 50.74 52.96 64.71 45.59 21.57
In-File 25.99 33.84 45.11 4.04 39.18 43.24 52.45 34.80 11.27
RepoCoder 46.22 50.31 57.23 15.53 48.39 51.01 60.78 46.08 20.59
Rambo 50.44 53.20 61.70 18.94 51.03 53.05 69.61 46.57 23.04
GPT3.5 - Turbo (175B) Rambo Oracle 51.21 53.91 60.85 19.15 67.8 67.96 76.96 61.76 35.29
RepoCoder Oracle 47.82 51.43 62.13 17.02 62.76 64.85 72.06 57.84 31.37
In-File 33.93 39.25 53.11 6.38 63.16 65.66 76.96 60.78 26.96
RepoCoder 44.41 49.64 58.09 15.74 63.52 65.39 74.02 59.80 31.86
Rambo 49.58 52.71 58.94 19.15 66.29 67.82 76.47 63.73 32.84
@@
Table II presents the MBC performance of Rambo compared to the state-of-the-art approaches using several representative code LLMs across our proposed benchmark and the Defects4J. As seen, Rambo exhibits significant improvements over In-File, RepoCoder, and even RepoCoder Oracle. Even when compared to its own Oracle (Rambo Oracle), Rambo maintains a highly competitive performance. For example, with GPT-3.5 Turbo on Rambo’s benchmark, Rambo achieves a significant increase of about 12–46% in BLEU and 6–34% in CodeBLEU compared to those of RepoCoder and In-File. Similarly, on the Defects4J dataset with Code Llama, Rambo shows consistent gains, outperforming the other approaches by 2–18% in BLEU and 3–17% in CodeBLEU. These results highlight Rambo’s ability to consistently generate syntactically accurate and high-quality code across diverse repositories, outperforming RepoCoder by a substantial margin in both BLEU and CodeBLEU metrics. This shows Rambo’s ability to generate syntactically and semantically accurate code with fewer errors and effectively leverage repo-specific code elements in generated bodies.
@@
Moreover, Rambo excels in generating method bodies that is compilable and exactly match the reference implementations, as shown by its significantly better CR (Compilation Rate) and Exact Match (EM) rates. With DeepSeek Coder on Rambo’s benchmark, Rambo’s EM rate improves 22% and 3X the corresponding figures of RepoCoder and In-File, respectively. With the same code LLM, Rambo delivers substantial CR improvements of 8% and 15% compared to RepoCoder in Defects4J. In terms of pass@k, Rambo also outperforms with 2/3 studied code LLMs, with the improvements of 1–33%, further affirming its effectiveness in generating functionally correct code.
@@
{mdframed}
[backgroundcolor=light-gray, skipabove=5pt, skipbelow=0pt, nobreak=false ]{mdtheorem}namelabel Answer to RQ1: Our results show Rambo’s consistent superiority over existing repo-level MBC approaches and its competitive performance compared to Rambo Oracle. Rambo not only outperforms current solutions but also sets a new benchmark by exceeding the anticipated upper-bound performance of RepoCoder, showing its robustness, accuracy, and practical utility in real-world code generation tasks.
@@
Fig. 2 shows an MBC problem in repository Spring Cloud Gateway [^1] https://github.com/spring-cloud/spring-cloud-gateway with the infilling method delete simplified in Fig. 2(a). The delete method in the InMemoryRouteDefinitionRepository class removes a RouteDefinition by its ID. It performs a Route Existence Check by verifying if the route exists in the map. If found, the method removes it and returns an empty Mono<Void> (Success Case). If the route ID is not present, it uses Mono.defer to lazily create and return a Mono.error emitting a NotFoundException (Error Handling). In this case, since the other declared methods in the repository are quite different from delete, the context provided to RepoCoder and its Oracle is insufficient. As shown in Fig 2(b), the body generated by RepoCoder and its Oracle completely omits both the Route Existence Check and Error Handling.
On the other hand, the sketch code generated by Rambo (Fig. 2(c)) includes the Route Existence Check and Success Case, but still misinterprets the error handling. Based on the sketch, Rambo identifies the essential elements and their usages within the repository. One relevant usage is the save method in the RedisRouteDefinitionRepository class (Fig. 2(d)). With this relevant usage as a reference where Mono.error is used as expected (lines 8–9 in save), Rambo and the code LLM are able to correctly generate the expected body, as shown in Fig. 2(e).

### v3

@@ 1 Introduction
We conducted several experiments to evaluate the performance of Rambo in repo-level method body completion using state-of-the-art code LLMs, including DeepSeek Coder [7], Code Llama [6], and GPT-3.5 Turbo [19]. These experiments were conducted across 40 Java projects with a total of 674 repo-level MBC problems. Our results demonstrate that Rambo significantly surpasses RepoCoder across all evaluated code LLMs and metrics. Specifically, Rambo achieves impressive improvements over the other MBC approaches in various performance metrics, by up to 46% in BLEU, up to 57% in CodeBLEU, up to 36% in Compilation Rate, and up to 3X in Exact Match rate. Remarkably, Rambo delivers up to 12% enhancement in Exact Match rate compared to the Oracle RAG-based method [15], which estimates the upper-bound of performance for RepoCoder and other methods employing the same pipeline. This substantial improvement shows Rambo’s superior capability in generating high-quality method bodies and sets a new benchmark in the task of repository-level method body completion.
@@
2. Advanced Retrieval Strategy for RAG-based MBC: We introduce a novel retrieval mechanism that identifies and utilizes essential repo-specific code elements, such as classes, methods, and fields, and their relevant usages to construct more accurate and contextually relevant contexts for RAG-based method body completion.
@@ 3.1 Repository-specific Context Retrieval
This section describes the retrieval phase, in Rambo to search for the repo-specific context relevant to the given infilling method, $C_{R}=\textit{retrieve}(m,R)$. Our retrieval approach consists of two phases, Essential Code Element Identification and Relevant Usage Extraction.
@@ 3.1.1 Essential Code Element Identification (EEI)
To enhance the relevance of the sketch, Rambo first searches for the set of methods, $L$, within the repository $R$ whose interface (input-output) and name are most similar to those of $m$. The set of methods $L$ serves as the reference points for understanding the expected structure and functionality of $m$. Then, Rambo employs a RAG-based method to generate a sketch code $\mathcal{S}$ of the method body using the retrieved methods $L$ and the surrounding context $C_{f}$. Although the sketch, generated as $\mathcal{S}=\mathscr{M}(m,L,C_{f})$, may not fully accurate, $\mathcal{S}$ includes potentially essential elements such as called methods, used fields, and types. These elements are then analyzed and extracted:
@@
where $\mathcal{M}_{m}$, $\mathcal{M}_{f}$, and $\mathcal{M}_{t}$ are the sets of called methods, accessed fields, and used types in the sketch $\mathcal{S}$, respectively. Note that analyzing the generated sketch code $\mathcal{S}$ by existing code analyzers (e.g., JDT or Joern [20]) is feasible because the code generated by pre-trained code models, especially code LLMs, is quite syntactically correct [21, 13]. However, the employed code LLM may hallucinate code, possibly generate the code elements not existing/declared in the repo, or use them inappropriately [14, 13].
To mitigate code LLM hallucination, after analyzing, Rambo searches for similar code elements already declared within the repository. This involves identifying the sets of methods, fields, and classes/types whose names are most similar to those found after analyzing $\mathcal{S}$:
@@
$E_{t}=\bigcup_{\begin{subarray}{c}e\in\mathcal{M}_{t}\end{subarray}}\left\{\arg\max\limits_{c\in\text{cls}(R)}\left(\text{name\_sim}(e,c)\right)\right\}$ (3)
@@ 3.1.2 Relevant Usage Extraction (RUE)
Once these usages are identified, they are ranked based on their relevancy to the functionality of the method $m$. The relevancy of a usage $u$ is calculated by function rel:
@@
where $\text{sim}(u,\mathcal{S})$ calculates the similarity between a usage $u\in U$ and the sketch $\mathcal{S}$ generated during EEI. The reason for this estimation of $\text{rel}(u)$ is that the body of $u$ expresses how an essential element $e$ is utilized and $\mathcal{S}$ roughly expresses how potentially essential elements should be used. Thus, if the usage of $e$ is also present in the sketch code, it indicates that the usage of $e$ in $u$ is likely relevant to $m$. This ranking ensures that the most relevant usages are prioritized, providing the code LLM with high-quality context that is both specific to the repository and relevant to the task at hand. Additionally, Rambo brings back the set of methods $L$ whose signature most closely matches that of $m$, serving as a structural and functional reference point for method $m$. The methods in $L$ and the top relevant usages are concatenated to form the repo-specific relevant context $C_{R}$. Finally, this context and the surrounding context $C_{f}$ of $m$ are fed into $\mathscr{M}$ to generate the final method body $\widehat{m_{b}}=\mathscr{M}(m,C_{f},C_{R})$.
@@ 3.2 Method Body Generation
To generate $m$’s body, we design a prompt template that seamlessly integrates the identified relevant context $C_{R}$ with the infilling method, embedding the context as code comments within the file (as illustrated in Fig. 1). The retrieved usages are arranged in ascending order of relevancy, with each code snippet annotated by its original file path. The number of snippets included in the prompt is optimized based on the available prompt length. This approach ensures that the prompt is packed with relevant information, facilitating the generation of a method body that is both functionally accurate and consistent with the repository’s coding standards.
@@ 4.1 Dataset
In this work, we evaluate the MBC performance of Rambo and other approaches on two large MBC benchmarks (Table 1). First, we utilize Defects4J [22] as in CoderUJB [23], a multi-task benchmark for several SE tasks such as functional code generation, test generation, defect detection, and automated program repair. In this work, we adapt the functional code generation task from Defects4J in CoderUJB to the MBC task, resulting in 204 MBC problems in 14 real open-source Java projects. Additionally, to evaluate the performance of the repo-level MBC approaches on more up-to-date repositories, we construct another benchmark which includes active high-quality open-source repositories from GitHub. Specifically, we collect 26 Java repositories (other than those in Defects4J) that meet the following criteria: they must have an open-source license, have been actively updated, possess over 2K stars, and be compilable. From each non-test class in the collected projects, we randomly select a method to generate an MBC problem and the corresponding reference, which is the method’s body. To ensure a reasonable difficulty level of the benchmark, we exclude all constructors, getters, setters, and too-short methods with fewer than 3 LOCs. As a result, we obtain a total of 470 MBC problems. In this work, for every MBC problem, we omitted the import section in the left context, $C_{f}^{\text{left}}$, to accurately reflect the coding activities in practice, that the necessary imports are usually unavailable before implementing the infilling method.
Table 1: Datasets statistics
@@ 4.2.1 Empirical Procedure
Baselines. We compared Rambo against the state-of-the-art repo-level MBC approaches:
@@
    RawRAG: This baseline employs the RAG strategy in its most straightforward form by retrieving relevant context solely based on the left context using a similarity search. The retrieved snippets are then directly fed into the code LLM for generation without further refinement of the retrieved content.
@@
    RLCoder [18]: RLCoder is a reinforcement learning framework that iteratively learns to retrieve high-quality context. In our work, RLCoder has been adapted to generate method bodies line by line until a complete body is formed.
@@
    RepoCoder [15]: RepoCoder employs a similarity-based retriever and a code LLM in an iterative retrieval-generation pipeline to generate complete unfinished code, e.g., unfinished lines, API invocations, or function/method bodies.
  5. 5.
    RepoCoder Oracle [15]: This method estimates the upper bound of RepoCoder’s performance by utilizing the ground truth code to search for relevant context before feeding into code LLMs in the RepoCoder. Note that this method performs a single retrieval process to obtain relevant code snippets.
  6. 6.
    Rambo Oracle: Inspired by RepoCoder Oracle [15], we estimated the upper bound of Rambo’s performance by its own Oracle. Within Rambo’s pipeline, this oracle method directly considers the ground truth code as the sketch code $\mathcal{s}$ in EEI and RUE. In this case, the essential elements are correctly identified, and their usages are ranked based on the similarities with the ground truth.
We used the implementation in their original papers for all the baseline approaches. For RepoCoder and RepoCoder Oracle, we used the hyper-parameters recommended in the original work [15]. Additionally, we applied RepoCoder with a single iteration at which RepoCoder achieves the best overall performance with our studied dataset. We evaluated the MBC performance of the approaches under study using several representative code LLMs, both open-source and commercial, including DeepSeek Coder 6.7B [7], Code Llama 7B [6], and GPT-3.5 Turbo [19] with about 175B parameters. For detailed implementation and other hyper-parameters, one can see our website [26]. Note that all experiments were run on a machine with an Intel Xeon Bronze 3104 (6 cores) @ 1.70GHz, 32GB RAM, and an NVDIA RTX A6000 GPU.
@@
  • Retrieval Technique: We evaluated how the retrieval technique powered by different embedding methods affects the MBC performance of Rambo.
@@
  • Relevant Context Size: We analyzed how changes in the size of the relevant context, $C_{R}$, influence the MBC performance.
  • Code LLM’s Size: We studied the effect of the size of the employed code LLM, $\mathscr{M}$, on Rambo’s performance, balancing computational efficiency and the quality of the generated method bodies.
@@
  • Repository Size: We examined the impact of repository sizes on Rambo’s ability to retrieve relevant context and generate method bodies.
@@ 4.2.2 Metrics
  • BLEU [27] is a widely used metric in NLP, assessing the similarity between generated and reference text by comparing overlapping n-grams. In the context of MBC, it measures surface-level similarity, focusing on the direct overlap of tokens between the generated and reference code.
  • CodeBLEU [28] extends the BLEU score by incorporating code-specific properties, such as syntax and semantic structure. It measures the alignment between the generated and the reference codes, capturing deeper structural elements.
  • Compilation Rate (CR) assesses the percentage of generated code bodies that successfully compile without errors. A higher compilation rate reflects the syntactical correctness of the generated code and its adherence to the language’s rules, making it a crucial metric for evaluating the practicality of code generation models.
  • pass@k [29] evaluates the generated code’s ability to pass a set of unit tests, measuring the percentage of solutions that pass all the given tests when considering the top-k predictions. A higher pass@k value indicates stronger model performance in producing functionally correct code. In this work, $k=1$ and pass@k is applicable for evaluation using Defects4J.
@@ 5.1 Answering RQ1: Accuracy and Comparison
Table 2: Repo-level MBC performance comparison on Rambo’s benchmark
Code LLM BLEU CodeBLEU CR EM
CodeLlama (7B) Rambo Oracle 41.83 46.22 51.28 14.89
RepoCoder Oracle 46.59 49.53 55.96 14.68
In-File 25.09 33.85 48.09 3.83
RawRAG 41.57 45.96 54.89 12.55
RLCoder 39.92 44.30 54.47 10.85
RepoCoder 42.98 47.66 55.32 12.77
Rambo 43.16 47.76 55.96 13.62
DeepSeek Coder (6.7B) Rambo Oracle 50.87 53.96 63.40 18.72
RepoCoder Oracle 50.04 52.48 58.51 17.02
In-File 25.99 33.84 45.11 4.04
RawRAG 46.92 50.97 60.21 15.32
RLCoder 45.08 48.58 55.96 14.26
RepoCoder 46.22 50.31 57.23 15.53
Rambo 50.44 53.20 61.70 18.94
GPT3.5 - Turbo (175B) Rambo Oracle 51.21 53.91 60.85 19.15
RepoCoder Oracle 47.82 51.43 62.13 17.02
In-File 33.93 39.25 53.11 6.38
RawRAG 42.90 45.27 55.53 12.55
RLCoder 45.12 49.48 58.72 15.32
RepoCoder 44.41 49.64 58.09 15.74
Rambo 49.58 52.71 58.94 19.15
Table 3: Repo-level MBC performance comparison on Defect4J dataset
Code LLM BLEU CodeBLEU CR pass@k EM
CodeLlama (7B) Rambo Oracle 54.19 55.90 64.71 46.08 19.12
RepoCoder Oracle 53.69 54.47 64.71 46.57 19.61
In-File 44.06 46.73 55.88 33.82 8.82
RawRAG 48.88 50.93 60.29 45.10 18.01
RLCoder 50.37 51.12 60.29 42.65 17.54
RepoCoder 51.27 53.46 60.78 46.57 20.10
Rambo 52.04 54.98 63.73 44.12 20.59
DeepSeek Coder (6.7B) Rambo Oracle 51.31 53.54 67.16 49.51 21.57
RepoCoder Oracle 50.74 52.96 64.71 45.59 21.57
In-File 39.18 43.24 52.45 34.80 11.27
RawRAG 44.96 48.47 61.27 45.10 18.01
RLCoder 45.40 48.10 57.84 45.59 18.01
RepoCoder 48.39 51.01 60.78 46.08 20.59
Rambo 51.03 53.05 69.61 46.57 23.04
GPT3.5 - Turbo (175B) Rambo Oracle 67.80 67.96 76.96 61.76 35.29
RepoCoder Oracle 62.76 64.85 72.06 57.84 31.37
In-File 63.16 65.66 76.96 60.78 26.96
RawRAG 66.09 67.36 75.00 61.27 31.28
RLCoder 65.78 66.67 75.98 61.76 32.84
RepoCoder 63.52 65.39 74.02 59.80 31.86
Rambo 66.29 67.82 76.47 63.73 32.84
@@
Table 2 and Table 3 present the MBC performance of Rambo compared to the state-of-the-art approaches using several representative code LLMs across our proposed benchmark and the Defects4J. As seen, Rambo exhibits significant improvements over In-File, RawRAG, RLCoder [18], RepoCoder [15], and even RepoCoder Oracle. Even when compared to its own Oracle (Rambo Oracle), Rambo maintains a highly competitive performance. For example, with GPT-3.5 Turbo on Rambo’s benchmark, Rambo achieves a significant increase of about 12–46% in BLEU and 6–34% in CodeBLEU compared to those of RepoCoder, RLCoder, RawRAG, and In-File. Similarly, on the Defects4J dataset with Code Llama, Rambo shows consistent gains, outperforming the other approaches by 2–18% in BLEU and 3–17% in CodeBLEU. These results highlight Rambo’s ability to consistently generate syntactically accurate and high-quality code across diverse repositories, outperforming RepoCoder and RLCoder by a substantial margin in both BLEU and CodeBLEU metrics. This shows Rambo’s ability to generate syntactically and semantically accurate code with fewer errors.
@@
Moreover, Rambo excels in generating method bodies that is compilable and exactly match the reference implementations, as shown by its significantly better CR (Compilation Rate) and Exact Match (EM) rates. With DeepSeek Coder on Rambo’s benchmark, Rambo’s EM rate improves 22% and 3X the corresponding figures of RepoCoder and In-File, respectively. With the same code LLM, Rambo delivers substantial CR improvements of 8%15% compared to RepoCoder in Defects4J. In terms of pass@k, Rambo also outperforms with 2/3 studied code LLMs, with the improvements of 1–33%, further affirming its effectiveness in generating functionally correct code.
@@
{mdtheorem} namelabel Answer to RQ1: Our results show Rambo’s consistent superiority over existing repo-level MBC approaches and its competitive performance compared to Rambo Oracle. Rambo not only outperforms existing approaches but also sets a new benchmark by exceeding the anticipated upper-bound performance of RepoCoder, showing its robustness, accuracy, and practical utility in real-world code generation tasks.
@@
Fig. 2 shows an MBC problem in repository Spring Cloud Gateway [^1] https://github.com/spring-cloud/spring-cloud-gateway with the infilling method delete simplified in Fig. 2(a). The delete method in the InMemoryRouteDefinition- Repository class removes a RouteDefinition by its ID. It performs a Route Existence Check by verifying if the route exists in the map. If found, the method removes it and returns an empty Mono<Void> (Success Case). If the route ID is not present, it uses Mono.defer to lazily create and return a Mono.error emitting a NotFoundException (Error Handling). In this case, since the other declared methods in the repository are quite different from delete, the context provided to RepoCoder and its Oracle is insufficient. As shown in Fig 2(b), the body generated by RepoCoder and its Oracle completely omits both the Route Existence Check and Error Handling.
On the other hand, the sketch code generated by Rambo (Fig. 2(c)) includes the Route Existence Check and Success Case, but still misinterprets the error handling. Based on the sketch, Rambo identifies the essential elements and their usages within the repository. One relevant usage is the save method in the RedisRouteDefinitionRepository class (Fig. 2(d)). With this relevant usage as a reference where Mono.error is used as expected (lines 8–9 in save), Rambo and the code LLM correctly generate the expected body for this MBC problem, as shown in Fig. 2(e).